OpenStudy (anonymous):

What's the solution? y > x-2 y <= (less than and equal to) -(1/3)x+3

OpenStudy (zzr0ck3r):

treat it like a equals sign but flip it if you ever multiply or divide by a negative number

OpenStudy (zzr0ck3r):

for the first one \(y>x-2\implies y\color{red}{+2}>x-2\color{red}{+2}\implies y+2>x\)

OpenStudy (anonymous):

Why did you add 2 to both sides?

OpenStudy (anonymous):

\[y >x-2\] \[y \le -\frac{ 1 }{ 3 }x +3\] What's the solution ^

OpenStudy (zzr0ck3r):

because you want to get x by itself. it is very much like this y=x-2 if we solve for x we add 2 to both sides.

OpenStudy (anonymous):

what do we do with the second equation then?

OpenStudy (zzr0ck3r):

\(y\le-\frac{1}{3}x+3\) subtract 3 from both sides \(y-3\le\frac{1}{3}x\) multiply both sides by 3 \(3(y-3)\ge x\)

OpenStudy (anonymous):

I need to find the solution of the system

OpenStudy (anonymous):

the system being those two inequalities

OpenStudy (anonymous):

>.< sorry if i'm being very confusing

OpenStudy (anonymous):

ohh i need to find the value of x c:

OpenStudy (anonymous):

Do you have a answer choice?

OpenStudy (anonymous):

I got y+2>x for the first I got y<=3-x/3 for the second

OpenStudy (anonymous):

3y-9=>x is what @zzr0ck3r wanted to say in his equation

OpenStudy (anonymous):

no answer choices

OpenStudy (texaschic101):

you do know that there is more then 1 solution

OpenStudy (anonymous):

So was I right?

OpenStudy (texaschic101):

I would probably graph each inequality and the solutions will be in the overlapping sections

OpenStudy (anonymous):

@texaschic101 I'm suppose to show it algebraically.

OpenStudy (blank ):

y>x-2 y<=-(1)/(3)*x+3 Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. x-2<y or y<=-(1)/(3)*x+3 Move all terms not containing x to the right-hand side of the inequality. x<y+2 or y<=-(1)/(3)*x+3 Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. x<y+2 or -(1)/(3)*x+3>=y Multiply -(1)/(3) by x to get -(x)/(3). x<y+2 or -(x)/(3)+3>=y Move all terms not containing x to the right-hand side of the inequality. x<y+2 or -(x)/(3)>=y-3 Multiply each term in the inequality by 3. x<y+2 or -x>=3y-9 Multiply -x by -1 to get x. x<y+2 or x<=3y*-1-9*-1 Divide each term in the inequality by -1. x<y+2 or x<=-3y+9 Hope this helps. This is the Union of the inequalities. You could also take a look at this: http://www.wolframalpha.com/input/?i=y%3Ex-2%2Cy%3C%3D-%281%2F3%29x%2B3